The last stage of a rocket, which is traveling at a speed of 8000 m/s, consist of two parts that are clamped together: a rocket case with a mass of 220.0 kg and a payload capsule with a mass of 120.0 kg. When the clamp is released, a compressed spring causes the two parts to separate with a relative speed of 910.0 m/s. (a) What are the speeds of the two parts after they have separated? Assume that all velocities are along the same line. m/s (rocket case) m/s (payload) (b) Find the total kinetic energy of the two parts before and after they separate. J (before) J (after) (c) Account for the difference.
Answer #1

Pi =P (220+ 120) (8000) = 220v + 1200, - and -V, = 910 a 12 = 910tv, (1) 2720000 = 2002007 2720000 = 220v, + 109200 + 12ou, *